Please complete the factored form of the polynomial
f(x)=6x^3+2x^2-16x+8

Answer by lwsshak3(11628)

Please complete the factored form of the polynomial
f(x)=6x^3+2x^2-16x+8
**
If you are allowed to use a calculator, you will find that 1 is one of the roots. The other two roots can be found by long division or synthetic division. I will use the latter method.
0).......6......2......-16.....8
1)...............6.........8....-8
..........6......8.......-8......0
f(x)=(x-1)(6x^2+8x-8)
6x^2+8x-8=0
3x^2+4x-4=0
(3x-2)(x+2)=0
3x-2=0
x=2/3
x+2=0
x=2
..
f(x)=(x-1)(3x-2)(x+2)